# Automated Testing Method for Text-to-Image Software

### v2

@@ 4. Experiment Design
Our evaluation was designed to answer the three main research questions in the experiments:
RQ1: Can ACTesting generate error-revealing tests for T2I software under the adaptability density-guided?
RQ2: Can three kinds of transformation operators effectively detect the defects of T2I software?
RQ3: How does the combination of different transformation operators impact the robustness of the tested T2I software?
@@ 4.1. Datasets and T2I Software
In this section, we introduce the dataset and software we implement in our experiments.
Dataset. We mainly use the MS-COCO (Caesar et al., 2018) as the seed sets, which is a widely recognized benchmark in the field of computer vision. It is introduced to address the need for rich annotations including labels of object class, object instances, object locations and the relationships to other objects. Importantly, the dataset stands out for its image captioning component, where each image is accompanied by at least five different captions provided by human annotators. This feature makes MS-COCO highly valuable at the intersection of vision and natural language processing.
T2I Software. To elucidate the effectiveness of the testing methods we propose, we implement four Text-to-Image software in the experiment. We employ the method of calling APIs to alleviate the potential bias and ensure the black-box nature of the experimental process. We briefly introduce this software as follows:
OpenJourney  [^2] https://www.openjourney.art/ As one of the most popular text-to-image software, OpenJourney is painting waves in the world of ai-generated art. It is often quoted as a free alternative to MidJourney as it is a Stable Diffusion model trained using thousands of Midjourney images from its v4 update.
Wan Xiang  [^3] https://wanxiang.aliyun.com/creation from Alibaba Cloud platform, is an evolving AI painting model that can create corresponding images or artworks from textual descriptions using machine learning and natural language processing technologies. This model is based on Alibaba’s proprietary combinatorial generative model, Composer (Huang et al., 2023).
Stable Diffusion XL (SD XL)  [^4] https://stablediffusionweb.com from Stability.ai platform, help users create descriptive images with shorter prompts and generate words within images. The model is a significant advancement in image generation capabilities, offering enhanced image composition and face generation that results in stunning visuals and realistic aesthetics.
DeepAI Image Generator  [^5] https://deepai.org/machine-learning-model/texT2Img from DeepAI platform, creates an image from scratch from a text description. It can be used to generate AI art or for general silliness. It provides the functions of AI image generation and AI image edition API call services for developers.
@@ 4.2. Experimental Setup
Implementing steps. We conduct testing experiments on 4062 seeds of text-image pairs from the MS-COCO validation dataset. We define 150 categories of entities and 50 kinds of relationships based on the distribution of the dataset. After that, we implement three adaptability density-guided operators (EC, ER_R, ER_A) to generate three mutation testing sets. To demonstrate the effectiveness of our testing methodology, we have selected a commonly used text mutation operator, random synonym substitution (SS), as our baseline. Furthermore, to validate the effectiveness of each of the three operators, we integrated EC with the other two operators (EC+ER_R, EC+ER_A), resulting in two types of compounded mutant test inputs. Lastly, we generate the synthetic images based on T2I software for all the test sets. In summary, we get 4 software * 4062 captions * 7 test sets = 113,736 synthetic images. All images are resized to 512 $\times$ 512. We calculate evaluation metrics based on all inputs and outputs and subsequently obtain the results.
Running environment. All experiments are performed on a Ubuntu 20.04.6 LTS server with RTX 3090 GPU. We implement ACTesting on Python 3.7. We experiment with our methods on four widely-used T2I software based on the public dataset MS-COCO (Caesar et al., 2018), and the targets of the testing process are black box software. We use the NLP Package Stanza  [^6] https://stanfordnlp.github.io/stanza/ powered by Stanford NLP Group and Bert (Devlin et al., 2018) to analyze the input text based on the Pytorch Framework.
@@ 4.3. Evaluation Metrics
Firstly, we follow the convention of the T2I model research community and employ several metrics. We adopt the Improved-Fréchet Inception Distance (I-FID), Improved-Inception Score (I-IS), and R-Precision (RP) based on (Bynagari, 2019)(Salimans et al., 2016)(Xu et al., 2018) and (Dinh et al., 2022) for evaluating the image realism and text relevance.
To compute the image realism metrics (I-FID and I-IS), we initially extract features using a pre-trained Inception-v3 network (Szegedy et al., 2016).To solve the overfitting problem, we adapt to calibrate the confidence score of the classifier (Inception-v3), to which we opt to apply the popular network calibration method of temperature scaling (Guo et al., 2017). The formula of I-FID is defined below.
@@
where $X_{r}\sim\mathcal{N}(\mu_{r},\Sigma_{r})$ and $X_{g}\sim\mathcal{N}(\mu_{g},\Sigma_{g})$ are the features of real images and generated images extracted by a pre-trained Inception-v3 model. Then, these two feature sets are represented as two multivariate Gaussian distributions. A lower I-FID value indicates superior image realism. The formula of I-IS is defined below.
@@
where $x$ is the generated image and $y_{c}$ is the class label. We consider that $p(y_{c}|x)$ must have low entropy and $p(y_{c})$ must have high entropy. Therefore the KL-divergence between $p(y_{c})|x)$ and $p(y_{c}|x)$ should be large. Higher IS value means better image quality and diversity.
The RP metric is widely utilized to assess the consistency between text and image. The principle behind RP involves re-querying a synthesized image using the initial input caption. Specifically, an image is generated based on a true textual description, amidst 99 other randomly chosen mismatched captions. This generated image is then used to search for the input description among 100 potential captions. The retrieval is considered successful if the image’s matching score with the original caption ranks the highest. For this multi-modal encoding process, we chose CLIP following (Dinh et al., 2022), a robust encoder for both text and images, which has been trained on an extensive dataset containing 400 million text-image pairs. A higher RP value indicates a better text-image matching degree.
Secondly, we define the error-detection ratio of entity ($Error_{e}$) and error ratio of relationship ($Error_{r}$) based on the MRs defined in Sec 3.2.
@@
Thirdly, we apply the accuracy referring to object detection and relationship retrieval for entities and relationships in generated images, detected by the scene graph generation model (Tang et al., 2020) (The model possesses a detection accuracy rate of 90%) to evaluate the text-image relevance. To more vividly demonstrate our testing method, we use 1 minus the accuracy rate to represent the miss-detection ratio. The two metrics are formulated as follows:
@@ 5.1. Answer to RQ1
We conduct our ACTesting on each software to explicitly demonstrate the effectiveness of our method. We also display the testing results of the metrics on real images, illustrating the differences between generative images and real images across various operators.
Results. Table 1 presents the I-FID, I-IS, and RP of all tested software on the seed sets and mutation testing sets. From the third column in Table 1, it’s evident that all four mutation operators effectively increased the I-FID values, thereby reducing the quality of the generated images. Among them, the ER_A operator showed an average increase of about 5%, ranking it first in effectiveness. Additionally, the I-IS values for all the software under test decrease on the four mutation test sets. The decreased range of EC, ER_A, and ER_R operators is 2.9% to 9.6%, 8% to 15%, and 1.7% to 8.8%. In the RP column, the three operators included in ACTesting significantly reduced the precision values of the seed test set, with the EC operator consistently decreasing the RP value by 16.4% to 21.1%. The ER_R and ER_A operators, on average, decreased by 7.5% and 8.4% respectively. In comparison, the performance of the SS operator was relatively mediocre.
Discussion. We discover that the degree of reduction in image quality and text-image relevance is closely related to transformation. The test sets mutated by the ER_R and ER_A operators achieve the higher I-FID values and the lower I-IS values. Additionally, the test set mutated by the EC operator leads to a consistent and substantial decrease in RP values. This is attributed to the EC operator disrupting the more common ER triples, compelling the software to generate more innovative results. As a consequence, this reduces the consistency between the text and the generated images. The reason for the lesser decline in the I-FID and I-IS metrics compared to RP is that our operators are primarily focused on detecting defects in text-image consistency. The three operators included in ACTesting all outperform the baseline SS, which serves as the baseline. The results indicate that ACTesting can generate error-revealing tests for T2I software.
Table 1. The I-FID, I-IS and RP of different software on the tests generated by four guidance approaches and original data
@@
OpenJourney ER_R 31.29 42.52 82.99%
ER_A 31.49 40.29 81.91%
SS 30.97 42.64 87.47%
MS COCO Real-Image 2.62 46.00 83.54%
@@ 5.2. Answer to RQ2
For each operator, ACTesting designs a related MR to detect the defect. As described in Section 3.2, we consider the degree of matching between images and text based on the ERs. Therefore, the corresponding MR for each operator will be categorized into two types: entities and relationships. These will be separately addressed in error reporting, and donating error rate of entities and relationships ($Error_{e}$ and $Error_{r}$). The MR for the baseline SS operator is set such that both entities and relationship sets remain unchanged.
Results. Figure.6 presents the $Error_{e}$ and $Error_{r}$ of different software on the tests generated by four operators (three proposed operators and one baseline). Subfigure a displays the error rate associated with the MR of entities. The error rate for the SS operator remains around 0.40, while the ER_R operator consistently stays around 0.60. The EC operator averages an error rate of 0.60, and the ER_A operator often ranks first, reaching its highest error rate of 0.69 in the OpenJourney software. Subfigure b illustrates the error rates for MR of relationships. In this case, the SS operator maintains error rates all below 0.4, while both the ER_R and EC operators exhibit error rates all above 0.6 but below 0.7. The ER_A operator ranks first, maintaining an error rate of 0.7 stably.
Discussion. Although the SS operator can also detect the defect within a certain range, our proposed operators can detect more erroneous behaviors of the tested software. The testing efficiency of the ER_A operator is around 1.75 that of the SS operator. From the experimental results of $Error_{e}$, ER_A is effective for most of the software both on entities and relationships. However, ER_R and EC show better performance when testing Wan Xiang. The results of $Error_{r}$ demonstrate that both of the three operators can detect the defects effectively and stably. Among the four software, Wan Xiang is more robust to the transformation. In conclusion, three kinds of transformation operators can effectively detect the defects of T2I software based on the related MR.
Figure 6. The error rates of different software on the tests generated by four mutation operators
Table 2. The ablation experiment results of different mutation operators.
@@ 5.3. Answer to RQ3
Table 2 shows the ablation experiment results of several mutation operators. To better demonstrate the effectiveness of three kinds of operators, we conduct the ablation experiments on the three operators, with the evaluation criteria being the miss detection rate for object detection and relationship retrieval in the generated images. We use Stanza to extract the ER from the input text as the ground truth. The metrics $Miss_{e}$ and $Miss_{r}$ we use are explained in Eq.19 and Eq.20.
@@
Results. Table 2 presents the ablation experiment results for six mutation test sets. The $Miss_{e}$ column values display the rates of missed detection for the four software under various mutation conditions. The table clearly shows that the tests conducted with the combined EC plus ER_R, and EC plus ER_A operators result in the highest rates of missed detection, surpassing those of the three individual operators. The value of $Miss_{e}$ for the EC+ER_A operator is, on average, 26% higher than that of the EC operator, and 29.1% higher than that of the ER_A operator. The average value of $Miss_{e}$ for the EC+ER_R operator exceeds that of the EC operator by 5.5% and surpasses the ER_R operator by a significant margin of 31.3%. The $Miss_{r}$ column reflects the rate of missed detections for relationships. It shows a trend consistent with that of the $Miss_{e}$ column. Specifically, the average $Miss_{r}$ value for the EC+ER_A operator exceeds those of the EC and ER_A operators by 5.1% and 7.2%, respectively. Although the EC+ER_R operator leads the EC operator on average $Miss_{r}$ by only 1.8%, it surpasses the ER_R operator by a notable margin of 10.8%, still demonstrating significant competitiveness.
@@
Discussion. The ablation study demonstrates that three kinds of operators are effective and flexible to use. Firstly, the value of $Miss_{e}$ and $Miss_{r}$ for each mutation operator is substantially higher than that of the seed test set, where EC+ER_A shows the best performance. This aligns with the trends observed in the common metrics presented in Table 1, further validating the effectiveness of this testing method. Secondly, we observe that the missed detection rate for entity recognition is higher than that for relationship recognition. This may be attributed to the higher precision of the object detection model and indicates that the accuracy of relationship recognition in image scene understanding still requires improvement. Additionally, the reason for the relatively lower competitiveness of the ER_R and EC+ER_R operators is that removing certain ER somewhat facilitates the T2I software in processing the remaining content. However, due to the removal of connective semantic properties, this operator still manages to detect defects in the software. This experiment also demonstrates that both Wan Xiang and Stable Diffusion XL exhibit commendable robustness.
5.4. Analysis
When we abstract the text-image pairs into triples, for text, it is not necessary to describe all relationships between entities, while for images, it is necessary to present all of them when constructing a visual scene. This difference in presentation formats can lead generative software to interpret the text as triples in different paradigms, leading to the omission of certain entities or relationships during the generation process.
Complex nested ERs often pose significant challenges for building models to generate images. In each input with nested ERs, these can be understood as multiple nodes and multiple edges. This means that the model needs to handle the joint probability distribution of each node and edge. Since nested relationships are typically highly correlated, the joint probability distribution between child nodes and parent nodes can influence each other, leading to a decrease in model performance and an inability to accurately predict node states.
Regarding the most commonly used diffusion models, which are based on modeling noise probability distributions, their loss function constrains the predicted noise to be close to the true noise at each step. When there are more or unfamiliar entity relationships and the joint probability distribution becomes more complex, the denoising process becomes more challenging. Their mechanisms tend to involve attention mechanisms, often selecting the parts they consider important for construction while discarding the parts that are difficult to build or that they consider unimportant.
5.5. Threats to Validity

### v3

@@ 4. Experiment Design
Our evaluation is designed to answer the three main research questions in the experiments:
RQ1: Can ACTesting generate low-quality output for T2I software?
RQ2: Can ACTesting generate error-revealing tests for T2I software?
RQ3: How does the combination of different mutation operators impact the generation quality of the tested T2I software?
@@ 4.1. Datasets and T2I Software
In this section, we introduce the dataset, software and baseline we used in our experiments.
Dataset. We use the MS-COCO (Caesar et al., 2018) as the seed sets, which is a widely recognized benchmark in the field of T2I. Importantly, the dataset stands out for its image captioning component, where each image is accompanied by at least five different captions provided by human annotators. These features make MS-COCO highly valuable at the intersection of vision and natural language processing.
T2I Software. To elucidate the effectiveness of the testing method we propose, we choose five widely-used T2I software using different models in the experiment. We briefly introduce the tested software as follows:
OpenJourney (PromptHero, 2023) is one of the most popular text-to-image software. OpenJourney is painting waves in ai-generated art. It is often quoted as a free alternative to MidJourney as it is a Stable Diffusion model trained using thousands of Midjourney images from its v4 update.
Wan Xiang (Alibaba, 2023) is an evolving Artificial Intelligence (AI) painting model  (Huang et al., 2023) that can create corresponding images or artworks from textual descriptions using machine learning and natural language processing technologies.
Stable Diffusion XL (Stability.ai, 2023) helps users create descriptive images with shorter prompts. The model is a significant advancement in image generation capabilities, offering enhanced image composition and face generation that results in stunning visuals and realistic aesthetics.
DeepAI Image Generator (DeepAI, 2023) creates an image from scratch basedon a text description. It can be used to generate AI art. It provides the functions of AI image generation API call services for developers.
DALLE2 (OpenAI, 2023) is an AI system that can create realistic images from a description in natural language. We choose DALLE2 because it is a more stable version of this series.
Baseline. ACTesting is the first black-box testing method for T2I software so there are no available baselines. To demonstrate the effectiveness of our testing methodology, we select a commonly used text mutation operator, random synonym substitution (SS), as our baseline. The reason is that SS does not bring unknown semantic errors in the text compared with swapping word positions or introducing spelling errors.
@@ 4.2. Experimental Setup
Implementing details. We conduct experiments on 4062 seeds of text-image pairs from the MS-COCO validation dataset. A total of six additional test sets were generated across all experiments based on the seeds. We test 150 categories of entities and 50 kinds of relationships based on the dataset setting and labels. In summary, we get 5 software * 4062 captions * 7 test sets = 142,170 synthetic images. All images are resized to 512 $\times$ 512. We calculate evaluation metrics based on all inputs and outputs and obtain the results. To the best of our knowledge, this is the largest experimental evaluation for T2I software.
Running environment. All experiments are performed on a Ubuntu 20.04.6 LTS server with two RTX 3090 GPUs. We implement ACTesting on Python 3.7 and Pytorch 1.10.
@@ 4.3. Evaluation Metrics
To concretize this measurement in the T2I task, we apply two metrics. Firstly, we adopt the I-FID, I-IS, and RP based on (Bynagari, 2019)(Salimans et al., 2016)(Xu et al., 2018) and (Dinh et al., 2022) for evaluating the image realism and text-image relevance. To compute the image realism metrics (I-FID and I-IS), we initially extract features using a pre-trained Inception-v3 network (Szegedy et al., 2016). To solve the overfitting problem, we adapt to calibrate the confidence score of the classifier (Inception-v3), to which we opt to apply the popular network calibration method of temperature scaling (Guo et al., 2017). The formula of I-FID is defined below:
@@
where $X_{r}\sim\mathcal{N}(\mu_{r},\Sigma_{r})$ and $X_{g}\sim\mathcal{N}(\mu_{g},\Sigma_{g})$ are the features of real images and generated images extracted by a pre-trained Inception-v3 model. Then, these two feature sets are represented as two multivariate Gaussian distributions. A lower I-FID value indicates superior image realism. The formula of I-IS is defined below:
@@
where $x$ is the generated image and $y_{c}$ is the class label. The KL-divergence between $p(y_{c})|x)$ and $p(y_{c}|x)$ should be large. Higher IS value means better image quality and diversity.
The RP metric is widely utilized to assess the consistency between text and image. The principle behind RP involves re-querying a synthesized image using the initial input caption. Specifically, an image is generated based on a true textual description amidst 99 other randomly chosen mismatched captions. This generated image is then used to search for the input description among 100 potential captions. The retrieval is considered successful if the image’s matching score with the original caption ranks the highest. A higher RP value indicates a better degree of text-image matching.
Secondly, we define the error-detection ratio of entity ($Error_{e}$) and relationship ($Error_{r}$) based on the MR specified in Sec 3.4.
@@
Thirdly, we apply the accuracy referring to object detection and relationship retrieval for both entities and relationships. We use the scene graph generation model (Tang et al., 2020) (The model possesses a detection accuracy rate of 90%) to detect the generated images. To more vividly demonstrate our testing method, we use 1 minus the accuracy rate to represent the miss-detection ratio. The two metrics are formulated as follows:
@@ 5.1. Answer to RQ1
Table 1. The I-FID, I-IS and RP of different software on the tests generated by five guidance approaches and original data.
@@
DALLE Orig 36.60 39.70 93.89%
SS 37.62 38.98 92.29%
ER_R 38.35 39.03 91.29%
ER_A 39.17 34.34 90.41%
EC 38.67 34.90 84.83%
MS COCO Real-image 2.62 51.25 95.54%
We conduct our ACTesting on each software to explicitly demonstrate its effectiveness. we implement three mutation operators (EC, ER_R, ER_A) and one baseline mutation operator (SS) to generate four mutation testing sets. We also display the testing results of the metrics on real images for comparison (Dinh et al., 2022). Results. Table 1 presents the I-FID, I-IS, and RP of all tested software on the seed sets and mutation testing sets. It can be seen that the three metrics of the baseline SS operator reflect a certain degree of decrease in generation image quality. The three mutation operators of ACTesting perform better. From the third column in Table 1, it’s evident that our operators effectively increased the I-FID values, thereby reducing the quality of the generated images. Among them, the ER_A operator showed an average increase of about 5%, ranking first in effectiveness. Additionally, the I-IS values for all the software under test also decrease. The decreased range of EC, ER_A, and ER_R operators is 2.9% to 9.6%, 8% to 15%, and 1.7% to 8.8%. In the RP column, the three mutation operators significantly reduced the precision values of the seed test set, with the EC operator consistently decreasing the RP value by 16.4% to 21.1%. The ER_R and ER_A operators, on average, decreased by 7.5% and 8.4% respectively.
Analysis. From the results, the three operators included in ACTesting all outperform the baseline SS. The test sets mutated by the ER_A operator achieve relatively higher I-FID values and lower I-IS values. This is because the ER_A operator introduces unfamiliar triplets and increases the difficulty and complexity of the generation process. Additionally, the test set mutated by the EC operator leads to a consistent and substantial decrease in RP values. This is attributed to the EC operator disrupting the natural ER triples, compelling the software to generate more innovative results. Consequently, this reduces the consistency between the text and the generated images. The reason for the lesser decline in the I-FID and I-IS metrics compared to RP is that our operators are primarily focused on detecting errors in text-image consistency. The results indicate that ACTesting can generate tests that reduce the quality of T2I software output.
@@ 5.2. Answer to RQ2
Figure 6. The error rates of different software on the tests generated by four mutation operators.
As described in Sec 3.4, we treat MR as the practical adequacy criteria, including definitions for entities and relationships. Therefore, the testing results are separately addressed in error reporting, calculated by the error rate of entities and relationships ($Error_{e}$ and $Error_{r}$). We define error detection of the SS operator when entities or relationships of the output are not exactly consistent with the input.
Results. Fig. 6 presents the detection error rates of five T2I software on the test inputs generated by four operators (three proposed operators and one baseline). The upper part shows the error rate of entities. The error rate for the SS operator remains around 0.40, while the ER_R and EC operators consistently stay around 0.60. The ER_A operator often ranks the highest, reaching its highest error rate of 0.69 in the OpenJourney software. The lower part illustrates the error rates for MR of relationships. In this case, the SS operator maintains error rates all below 0.4, while both the ER_R and EC operators exhibit error rates all above 0.6 but below 0.7. The ER_A operator ranks highest, maintaining an error rate of 0.7.
Analysis. Although the SS operator can also detect errors within a certain range, our proposed operators can detect more erroneous behaviors of the tested software. The testing efficiency of the ER_A operator is around 1.75 that of the SS operator. The reason is that ACTesting transforms entities and relationships, which allows the new text to encompass new testing scenarios. From the experimental results of $Error_{e}$, ER_A is effective for most software on entities and relationships. The results of $Error_{r}$ demonstrate that both of the three operators can detect the errors effectively and stably. In addition, By comparing with the results in Table 1, it is evident that the differences from the baseline reflected by the $Error_{e}$ and $Error_{r}$ are more pronounced. This is because the two metrics more accurately indicate the errors present in the generated images (the specific types of errors are described in Sec 6.2).
Table 2. The ablation experiment results of different mutation operators. Bold represents the highest missed detection rate, and underline represents the second-ranking.
@@ 5.3. Answer to RQ3
To better demonstrate the effectiveness of the three kinds of operators, we conduct ablation experiments. The evaluation criteria are the miss-detection rate for entities and relationships in the generated images. Due to the conflict between ER_R and ER_A operators, we combine the EC operator with the ER_R operator and ER_A operator, respectively. Table 2 shows the ablation experiment results of five T2I software.
@@
Results. The $Miss_{e}$ column displays the miss-detection rate of entities under various mutation conditions. The table clearly shows that the tests conducted with the combined EC plus ER_R, and EC plus ER_A operators result in the highest missed detection rates, surpassing those of the three individual operators. The value of $Miss_{e}$ for the EC+ER_A operator is, on average, 26% higher than that of the EC operator and 29.1% higher than that of the ER_A operator. The average value of $Miss_{e}$ for the EC+ER_R operator exceeds that of the EC operator by 5.5% and surpasses the ER_R operator by a significant margin of 31.3%. The $Miss_{r}$ column reflects the rate of missed detections for relationships. It shows a trend consistent with that of the $Miss_{e}$ column. Specifically, the average $Miss_{r}$ value for the EC+ER_A operator exceeds those of the EC and ER_A operators by 5.1% and 7.2%, respectively. Although the EC+ER_R operator leads the EC operator on average $Miss_{r}$ by only 1.8%, it surpasses the ER_R operator by a notable margin of 10.8%, still demonstrating significant competitiveness.
@@
Analysis. The ablation study demonstrates that three kinds of operators are effective and flexible to use. Firstly, the value of $Miss_{e}$ and $Miss_{r}$ for each mutation operator is substantially higher than that of the seed test set, where EC+ER_A shows the best performance. This aligns with the trends observed in the common metrics presented in Table 1, further validating the effectiveness of this testing method. Secondly, we observe that the missed detection rate for entity is higher than that for relationship. This may be attributed to the higher precision of the object detection model and indicates that the accuracy of relationship recognition in image scene understanding still requires improvement.
6. Discussion
6.1. False Positive and Negative
ACTesting suffers little from false positives and negatives. Due to the large number of experimental results, we randomly select 100 images for manual review. Although ACTesting utilizes high-precision detection and recognition models to calculate and analyze the testing results, we still encounter low false positive and false negative rates. Due to the strict setting of the MR, most of the false positive errors depend on the accuracy of the detection model; that is, the recognition model may mistake the background for the focal entity (the proportion is lower than 10%). The majority of false negatives of ACTesting come from the recognition model’s perception that the entity is partly distorted (Image 4 in Fig. 3), making it unable to be identified as a focal entity. Therefore, the false negative depends on the threshold of the detection models (8% in our setting).
6.2. Types of Detected Errors
(a) Entity Omission Error. The entities highlighted in red are absent in the images.
(b) Entity Distortion Error. The distortion part is outlined in red in the images)
(c) Relationship Construction Error. The relationships marked in green are incorrectly represented in the images.
Figure 7. Examples of each type of error reported by ACTesting. Images are generated by Stable Diffusion.
To further illustrate the errors detected by ACTesting, We provide two examples for each of the three error types, along with an indication of the reasons for the errors in Fig. 7. Firstly, the most prevalent error type is Entity Omission Error, commonly occurring when ACTesting employs the EC operator. As illustrated in Fig. 7 (a), the entity “child” is replaced with “surfboard” and “rock” is replaced with “guy” in the text; however, the generated images omit these entities. Secondly, while the T2I software is capable of generating entities, the quality of the generation is unsatisfactory. Fig. 7 (b) presents examples of Entity Distortion Error, which can be tested by all mutation operators. These two types of errors are related to entities. Both missing and distorted image entities are classified as entity generation failures by the object detection model, contributing to $Error_{e}$ and $Miss_{e}$. The third error type is Relationship Construction Error, which indicates incorrect relationship representation between entities in the images (e.g., the grey cat is “under” the umbrella rather than “on top of” it in Fig. 7 (c)), contributing to $Error_{r}$ and $Miss_{r}$. It is notable that (1) the three types of error can occur at the same time, as shown in the first image in Fig. 7 (c); (2) If an entity is classified as a generation failure, then any relationships associated with that entity will also be deemed failures.
6.3. Analysis and Insight
ER relationships can be incorporated into the construction and training of generative models. When we abstract the text-image pairs into ER triples, for text, it is not necessary to describe all relationships between entities, while for images, it is necessary to present all of them when constructing a visual scene. This difference in presentation formats can lead generative software to interpret the text as triples in different paradigms, leading to omission during the generation process. Complex or unreal-world ERs in the input text can pose significant challenges for T2I model training on natural scenes. Regarding the most commonly used diffusion models, which are based on modeling noise probability distributions, their loss function constrains the predicted noise to be close to the true noise at each step. When there are more or unfamiliar elements and the joint probability distribution becomes more complex, the denoising process becomes more challenging. Therefore, if ER relationships can be used as guided input during the model construction phase, it may enhance the model’s ability to construct the core parts of an image.
7. Threats to Validity